# Orthogonal projections of hypercubes

### v2

@@ VII.5 The hypercubic energy landscape of the three-ring system
VII.6 Fraction of explained variance
We investigate the fraction of explained variance by the PCs of the hypercubic energy landscape in Fig. 11. to see how the frustration of the system influences the PCs. For the one-ring system [Fig. 11(a1)], PC1 dominates the explained variance by the PCs, which is consistent with PC1 being proportional to the ground state. Turning to the two-ring and three-ring systems [Figs. 11(a2) and 11(a3)], the first PCs explain less than 90% of the variance. This decrease in the explained variance by the first component arises from the degeneracy of the ground states. PC2 explains the variance of degenerate ground states. For the three-ring system [Fig. 11(a3)], the first three PCs are required to explain more than 80% of the variance, and the second and third fractions of explained variance are the same, indicating that PC2 and PC3 share the regularity: the symmetry of the system.
Figure 12: Same as Figs. 10(a3), 10(b3), 10(c3), and 10(d3) but by different PCs. (a) The hypercubic energy landscape of the three-ring system by PC1 and PC3. (b) PC1 loading of (a) and the color of the bar matches the color of the corresponding biplot basis in (a). (c) Same as (b) but for PC3. (d) The Ising spin interaction networks of the three-ring systems, where the nodes are colored by the angle of the corresponding biplot basis in (a). (e–h) Same as (a–d) but by PC2 and PC3. Notice that in (a) and (e), several biplot bases are overlapped.
@@ VII.7 The hypercubic energy landscape of the three-ring system through PC3
The fraction of explained variance of the three-ring system [Fig. 11(a3)] leads us to visualize the hypercubic energy landscapes of three-ring101010 See the Supplemental Material [105] for hypercubic energy landscapes of two-ring systems using PC3. systems using PC3. In Fig. 12(a), the hypercubic energy landscape of three-ring systems, the same as Fig. 10(a3), is shown but by PC1 and PC3. We find that with PC1 and PC3, unlike the projection by PC1 and PC2, the shape of the hypercubic energy landscape looks similar to a parallelogram, not a hexagon. The biplot bases on the top right of Fig. 12(a) are only six arrows despite the system having 15 spins, indicating numerous overlaps of the biplot bases. To check which spins are overlapped, we show the loading of PC1 and PC3 in Figs. 12(b) and 12(c), and the interaction network in Fig. 12(d), coloring them by the angle of the biplot basis. PC3 loading [Fig. 12(c)], in particular, has a nonuniform magnitude of the elements; the magnitude of the elements of spins 1, 2, 5, 6, 8, and 9 are larger than the others. Those spins belong to the lower left hexagon of the interaction network in Fig. 12(d). The interaction network colored by the angle of the biplot basis in Fig. 12(d) illustrates how the biplot bases captures the correlation of spins. As indicated by PC3 loading [Fig. 12(c)], the spins on the lower left hexagon (spins 1, 2, 5, 6, 8, and 9) form the (anti)correlated group. The rest of the two rings form a large (anti)correlated group. PC3 loading creates the hypercubic energy landscape, emphasizing the lower left hexagon of the interaction network.
We further investigate the hypercubic energy landscape of the three-ring system by PC2 and PC3 to understand the symmetry of the system. The hypercubic energy landscape of the three-ring system by PC2 and PC3 [Fig. 12(e)], looks like a hexagon, similar to the projection by PC1 and PC2. Nevertheless, the landscape is more symmetric and regular than that by PC1 and PC2. Around the origin, the connected six ground states also form a regular hexagon, but they are overlapping with the other group of six ground states, cf. Fig. 10(a3). The angles of the biplot bases are almost uniform (with some overlapping), and the lengths of them are nearly the same, resulting in a symmetric and regular hypercubic energy landscape; this projection is the same kind as the isometric projection, but that of a nine-dimensional hypercube.
@@ VIII Conclusions
Our contributions are summarized as follows: (1) We introduce the Hamming and fractal projection of hypercubes in Sec. II. (2) We propose interpreting the biplot bases of PCA as the bases of the orthogonal projection of hypercubes in Sec. III. (3) We find that the PC1 loading of PCA on hypercubes is equivalent to the weighted hypercubic vertices in Sec. IV, and validate it analytically and numerically in Sec. V. (4) We reveal numerically and analytically that the quality of the orthogonal projection of hypercubes tends to be worse around the origin of the projected space in Sec. VI. (5) We visualize the hypercubic energy landscapes of finite artificial spin-ice systems and extract physical interpretations of them in Sec. VII.
@@ B.4 Minimum inner-product error formulation
where $\kappa_{i}$ is a nonzero coefficient $\kappa_{i}\neq 0$ and $\sum_{i=1}^{n}\kappa_{i}^{2}=1$. With this expansion, $\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}\lambda_{i}\acute{\bm{u}}_{i}$, $\bm{u}^{\top}\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}^{2}\lambda_{i}$, and $\acute{\bm{\varSigma}}\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}\lambda_{i}^{2}\acute{\bm{u}}_{i}$. Plugging these into Eq. (79), we obtain
@@
$=$ $\frac{1}{2}\left[\operatorname{tr}\left(\acute{\bm{\varSigma}}\right)-\lambda_{1}^{2}\right],$ (87)
@@ D.1 General bound of projected coordinates
In general, the orthogonally projected coordinates of hypercubic vertices are bounded by a circle with radius $\sqrt{N}$. This is because we have $\left|\bm{r}\left(\bm{s}\right)\right|\leq\sqrt{N}$ for any vertex. To show this, we consider the upper bound of $\left|\bm{r}\right|^{2}$,

### v3

@@ VII.7 The hypercubic energy landscape of the three-ring system through PC3
The fraction of explained variance of the three-ring system [Fig. 11(a3)] leads us to visualize the hypercubic energy landscapes of three-ring101010 See the Supplemental Material [109] for hypercubic energy landscapes of two-ring systems using PC3. systems using PC3. In Fig. 12(a), the hypercubic energy landscape of three-ring systems, the same as Fig. 10(a3), is shown but by PC1 and PC3. We find that with PC1 and PC3, unlike the projection by PC1 and PC2, the shape of the hypercubic energy landscape looks similar to a parallelogram, not a hexagon. The biplot basis on the top right of Fig. 12(a) are only six arrows despite the system having 15 spins, indicating numerous overlaps of the biplot basis. To check which spins are overlapped, we show the loading of PC1 and PC3 in Figs. 12(b) and 12(c), and the interaction network in Fig. 12(d), coloring them by the angle of the biplot basis. PC3 loading [Fig. 12(c)], in particular, has a nonuniform magnitude of the elements; the magnitude of the elements of spins 1, 2, 5, 6, 8, and 9 are larger than the others. Those spins belong to the lower left hexagon of the interaction network in Fig. 12(d). The interaction network colored by the angle of the biplot basis in Fig. 12(d) illustrates how the biplot basis capture the correlation of spins. As indicated by PC3 loading [Fig. 12(c)], the spins on the lower left hexagon (spins 1, 2, 5, 6, 8, and 9) form the (anti)correlated group. The rest of the two rings form a large (anti)correlated group. PC3 loading creates the hypercubic energy landscape, emphasizing the lower left hexagon of the interaction network.
Figure 12: Same as Figs. 10(a3), 10(b3), 10(c3), and 10(d3) but by different PCs. (a) The hypercubic energy landscape of the three-ring system by PC1 and PC3. (b) PC1 loading of (a) and the color of the bar matches the color of the corresponding biplot basis in (a). (c) Same as (b) but for PC3. (d) The Ising spin interaction networks of the three-ring systems, where the nodes are colored by the angle of the corresponding biplot basis in (a). (e–h) Same as (a–d) but by PC2 and PC3. Notice that in (a) and (e), several biplot basis are overlapped.
We further investigate the hypercubic energy landscape of the three-ring system by PC2 and PC3 to understand the symmetry of the system. The hypercubic energy landscape of the three-ring system by PC2 and PC3 [Fig. 12(e)], looks like a hexagon, similar to the projection by PC1 and PC2. Nevertheless, the landscape is more symmetric and regular than that by PC1 and PC2. Around the origin, the connected six ground states also form a regular hexagon, but they are overlapping with the other group of six ground states, cf. Fig. 10(a3). The angles of the biplot basis are almost uniform (with some overlapping), and the lengths of them are nearly the same, resulting in a symmetric and regular hypercubic energy landscape; this projection is the same kind as the isometric projection, but that of a nine-dimensional hypercube.
@@ VIII Conclusions
Our contributions are summarized as follows: (1) We introduce the Hamming and fractal projection of hypercubes in Sec. II. (2) We propose interpreting the biplot basis of PCA as the basis of the orthogonal projection of hypercubes in Sec. III. (3) We find that the PC1 loading of PCA on hypercubes is equivalent to the weighted hypercubic vertices in Sec. IV, and validate it analytically and numerically in Sec. V. (4) We reveal numerically and analytically that the quality of the orthogonal projection of hypercubes tends to be worse around the origin of the projected space in Sec. VI. (5) We visualize the hypercubic energy landscapes of finite artificial spin-ice systems and extract physical interpretations of them in Sec. VII.
@@ B.4 Minimum inner-product error formulation
where $\kappa_{i}$ is a coefficient satisfying $\kappa_{i}^{2}\in\left[0,1\right)$ and $\sum_{i=1}^{n}\kappa_{i}^{2}=1$. With this expansion, $\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}\lambda_{i}\acute{\bm{u}}_{i}$, $\bm{u}^{\top}\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}^{2}\lambda_{i}$, and $\acute{\bm{\varSigma}}\acute{\bm{\varSigma}}\bm{u}=\sum_{i=1}^{n}\kappa_{i}\lambda_{i}^{2}\acute{\bm{u}}_{i}$. Plugging these into Eq. (79), we obtain
@@
$=$ $\frac{1}{2}\left[\operatorname{tr}\left(\acute{\bm{\varSigma}}^{2}\right)-\lambda_{1}^{2}\right],$ (87)
@@ D.1 General bound of projected coordinates
In general, the orthogonally projected coordinates of hypercubic vertices are bounded by a circle with radius $\sqrt{N}$. This is because we have $\left|\bm{r}\left(\bm{s}\right)\right|\leq\sqrt{N}$ [Eq. (44)] for any vertex. To show this, we consider the upper bound of $\left|\bm{r}\right|^{2}$,